the energy used by an electric iron for 1 minute is 33 kJ, at a voltage of 220 volts. How large the current is in the electric iron?



1
Expert's answer
2022-02-16T13:24:27-0500

Given:

"t=60\:\rm s"

"W=33000\:\rm J"

"V=220\;\rm V"


The work done by electric field

"W=IVt"

Thus, the electric current

"I=\frac{W}{Vt}=\frac{33000}{220*60}=2.5\: \rm A"